Unit 1 Review:


3) Using the definition of an inverse, explain why f(x) and g(x) inverses of one another?

Why or why not?


1
Expert's answer
2021-09-20T16:29:50-0400

They are not inverses of one another. For them to be, then "f(g(x))=g(f(x))=x"


But "f(x)=3^{2x} +1 \text{ and } g(x)=\log{(x-1)}\\
f(g(x))=x^2-2x+2\neq x\\
g(f(x))=2x\neq x\\
\text{Hence they are not inverses of one another}"